This is an macroeconomics question:

Suppose that at first the price of a

bag of coffee

is

$5

and the price of a

box of tea

is

$3

.

Then, the price of a

bag of coffee

changes to

$10

and the price of a

box of tea

changes to

$7

.

What has happened the money prices and relative prices of these two goods?

A.

The money and relative prices of a

bag of coffee

and a

box of tea

have not changed.

B.

The money price of a

bag of coffee

and a

box of tea

have fallen and the relative price of a

box of tea

and a

bag of coffee

have risen.

C.

The money price of a

bag of coffee

and a

box of tea

have risen and the relative prices of a

box of tea

and a

bag of coffee

have fallen.

D.

The money price of a

bag of coffee

and a

box of tea

have risen and the relative price of a

bag of coffee

has fallen while relative price of a

box of tea

has risen.
Answer #1

Relative price of coffee initially = 5/3= 1.67

New relative of price coffee = 10/7= 1.43

So D is correct